OpenStudy (anonymous):

The volume of the solid formed by revolving the region bounded by y = 1/x, y = x and x = 2 about the x-axis. I need some help setting up the formula for the washer method. I think it is pi([2]^2 - [(1/x) + x^2]^2) dx.

OpenStudy (dumbcow):

 washer method: \[V = \pi \int\limits_a^b R^2 - r^2\] \[V = \pi \int\limits_1^2 x^2 - \frac{1}{x^2}\]